OpenStudy (anonymous):

find the extreme values using Lagrange multiplier subject to the given constraints f(x,y,z) = x + 2y, x + y + z = 1 y^2 + z^2 = 4

OpenStudy (bahrom7893):

I used to be good at this at one point. Let me see if I remember anything.

OpenStudy (anonymous):

okay, i know how to do them. i'm just having some trouble with this one

OpenStudy (bahrom7893):

i'll cal lambda L fx = L*gx fy=L*gy

OpenStudy (bahrom7893):

hold on i dont even remember what to do next lol, getting my notes.

OpenStudy (bahrom7893):

Okay: fx = L*gx fy = L*gy fz = L*gz So: 1=L*1 .. hmm i'm lost why do u have two constraints. Zarkon can u help us out?

OpenStudy (zarkon):

\[g(x,y,z)=x+y+z-1\] \[h(x,y,z)=y^2+z^2-4\] you want \[\nabla f=\lambda\nabla g+\gamma\nabla h\]

OpenStudy (zarkon):

do you see how to proceed?

OpenStudy (anonymous):

is that lambda h?

OpenStudy (zarkon):

I used gamma...but you could use lambda1 and lambda 2

OpenStudy (anonymous):

so i'm going to have 4 unknowns? can you solve it so i can see how to do it?

OpenStudy (zarkon):

I have a solution...you show me what you can do and I'll let you know if it is correct.

OpenStudy (zarkon):

using \[\nabla f=\lambda_1\nabla g+\lambda_2\nabla h\] can you find \[f_x=\lambda_1g_x+\lambda_2 h_x\]

OpenStudy (anonymous):

okay so i have (0,2,0) = \[\lambda(1,1,1) + \beta(0,2y,2z)\]

OpenStudy (zarkon):

why (0,2,0)

OpenStudy (anonymous):

(1,2,0)

OpenStudy (zarkon):

ok

OpenStudy (zarkon):

so what is \(\lambda\)

OpenStudy (anonymous):

1, 2-2ybeta,-2zbeta

OpenStudy (zarkon):

the first equation gives us \[1=1\cdot\lambda=0\cdot\beta\] thus \[\lambda=1\]

OpenStudy (anonymous):

yup got that

OpenStudy (zarkon):

\[1=1\cdot\lambda+0\cdot\beta\]

OpenStudy (zarkon):

ok...solve the other 2 equations for beta...then set them equal to each other

OpenStudy (anonymous):

should i ignore lambda?

OpenStudy (zarkon):

lambda=1

OpenStudy (anonymous):

oh yeah!

OpenStudy (zarkon):

you should get a relationship between y and z

OpenStudy (anonymous):

b = 1/2y and b = -1/2z

OpenStudy (anonymous):

so i set those equal

OpenStudy (anonymous):

y = -z

OpenStudy (zarkon):

yes...since they are both equal to beta

OpenStudy (zarkon):

correct

OpenStudy (zarkon):

nopw look at the equation \[y^2+z^2=4\]

OpenStudy (anonymous):

z = +/- root 2

OpenStudy (zarkon):

yes

OpenStudy (zarkon):

so what is y then?

OpenStudy (anonymous):

so y is the same, well -/+ root 2

OpenStudy (anonymous):

just the opposite

OpenStudy (zarkon):

correct...opposite sign

OpenStudy (zarkon):

now look at x+y+z=1

OpenStudy (anonymous):

and x = 1!

OpenStudy (zarkon):

yep

OpenStudy (anonymous):

awesome thank you so much!

OpenStudy (zarkon):

np